OpenStudy (anonymous):

solve for t: t/4=4/8

jimthompson5910 (jim_thompson5910):

cross multiply to get 8t = 4*4 8t = 16 I'll let you finish

OpenStudy (anonymous):

Do the same thing we did in the last one! :D Now your variable is t instead of x.

OpenStudy (anonymous):

okay let me try this 1 sec

OpenStudy (anonymous):

t=2?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

yay!! but would i do the same thing for this one? 18/20=45/t? solviing for t again but cross simplifying doesnt go in evenly from what i got..

jimthompson5910 (jim_thompson5910):

18/20=45/t 18t=20*45 18t = 900 keep going...

OpenStudy (anonymous):

18 would become 18t

OpenStudy (anonymous):

ok 1 sec

OpenStudy (anonymous):

divide 18 to both sides?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

t=50:)

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (anonymous):

thankks!!